if a =1, b =2, c= -3 find the value of c-ab
c+ab

Answer by elima(1433)

You are given the values of a,b & c; so just plug those into the equations to solve;
a =1, b =2, c= -3 find the value of c-ab
c+ab
c-ab
-3-1(2)
-3-2=-5
c+ab
-3+1(2)=-3+2=-1
=)